# Variational Diffusion Posterior Sampling with Midpoint Guidance

### v1

@@ 3 The MGPS algorithm
First, consider the DDPM forward process $q_{k|j}(x_{k}|x_{j})\vcentcolon=\mathcal{N}(x_{k};\sqrt{\alpha_{k}/\alpha_{j}}x_{j},(1-\alpha_{k}/\alpha_{j})\mathrm{I}_{d})$ defined for all $(j,k)\in\llbracket 0,n\rrbracket^{2}$ such that $j<k$. Define also the joint distribution obtained by initializing the same process with the posterior of interest (2.1)
$\pi_{0:n}(x_{0:n})\vcentcolon=\pi(x_{0})\prod_{k=0}^{n-1}q_{k+1|k}(x_{k+1}|x_{k})\,.$ (3.1)
Note that the identity $\smash{q_{k|0}(x_{k}|x_{0})=\int\prod_{j=0}^{k-1}q_{j+1|j}(x_{j+1}|x_{j})\,\mathrm{d}x_{1:k-1}}$ implies that the $k$-th marginal of this distribution is $\pi_{k}$ as defined in (3.1). The time-reversed decomposition of (3.1) writes
$\pi_{0:n}(x_{0:n})\vcentcolon=\pi_{n}(x_{n})\prod_{k=0}^{n-1}\smash{\pi}_{k|k+1}(x_{k}|x_{k+1})\,,$ (3.2)
$\smash{\pi}_{i|j}(x_{i}|x_{j})=\pi_{i}(x_{i})q_{j|i}(x_{j}|x_{i})\big{/}\pi_{j}(x_{j})\,.$ (3.3)
Note that following (2.2), we have that $\smash{\pi}_{i|j}(x_{i}|x_{j})\propto g_{i}(x_{i})q_{i|j}(x_{i}|x_{j})$. From this decomposition an exact sample from (2.1) is obtained by first sampling $X_{n}$ from $\pi_{n}$ and then recursively sampling $\smash{X_{k}\sim\smash{\pi}_{k|k+1}(\cdot|X_{k+1})}$ for $k\in\llbracket 0,n-1\rrbracket$. However, the backward-transition kernels (3.3) are difficult to approximate, as they also require knowledge of the intractable potentials $g_{k}$ according to (2.2). To mitigate the effect of the introduced approximation error associated with these potentials, we consider a new approach to the approximate sampling from $\smash{\pi}_{k|k+1}$ based on the decomposition (3.4) given in Lemma 3.1. First, define, for $\ell\in\llbracket 1,k\rrbracket$, the bridge kernel $q_{k|\ell,k+1}(x_{k}|x_{\ell},x_{k+1})\propto q_{k|\ell}(x_{k}|x_{\ell})q_{k+1|k}(x_{k+1}|x_{k})$, which is the density of a Gaussian distribution; see Section A.1. By convention we assume that $q_{k|k,k+1}(\cdot|x_{k},x_{k+1})=\delta_{x_{k}}$.
@@
$\smash{\pi}_{k|k+1}(x_{k}|x_{k+1})=\int q_{k|{\ell_{k}},k+1}(x_{k}|x_{\ell_{k}},x_{k+1})\smash{\pi}_{{\ell_{k}}|k+1}(x_{\ell_{k}}|x_{k+1})\,\mathrm{d}x_{\ell_{k}}\,.$ (3.4)
The proof is provided in Section A.3. From an approximate inference standpoint, this alternative decomposition is appealing for sampling from $\smash{\pi}_{k|k+1}$. First, notice that sampling from the bridge kernel $q_{k|{\ell_{k}},k+1}$ is straightforward as it involve sampling from a Gaussian distribution, hence we only need to focus our efforts on approximate sampling from $\smash{\smash{\pi}_{{\ell_{k}}|k+1}}$. This observation is the basis of the introduction of the following surrogate model approximating $\pi_{0:n}$ that we propose. For all $k\in\llbracket 1,n-1\rrbracket$ and ${\ell_{k}}\in\llbracket 0,k\rrbracket$, define the approximation
$\smash{\hat{\pi}^{\ell}_{k|k+1}(x_{k}|x_{k+1})}\vcentcolon=\int q_{k|{\ell_{k}},k+1}(x_{k}|x_{\ell_{k}},x_{k+1})\smash{\hat{\pi}^{\theta}_{{\ell_{k}}|k+1}(x_{\ell_{k}}|x_{k+1})}\,\mathrm{d}x_{{\ell_{k}}}$ (3.5)
of (3.4), where
$\smash{\hat{\pi}^{\theta}_{{\ell_{k}}|k+1}(x_{\ell_{k}}|x_{k+1})}\propto\hat{g}^{\theta}_{{\ell_{k}}}(x_{\ell_{k}})p^{\theta}_{{\ell_{k}}|k+1}(x_{\ell_{k}}|x_{k+1})\,,\quad\hat{g}^{\theta}_{{\ell_{k}}}(x_{\ell_{k}})\vcentcolon=g(\smash{m^{\theta}_{0|{\ell_{k}}}}(x_{\ell_{k}}))\,,$ (3.6)
and $\smash{p^{\theta}_{{\ell_{k}}|k+1}(\cdot|x_{k+1})\vcentcolon=q_{{\ell_{k}}|0,k+1}(\cdot|\smash{m^{\theta}_{0|k+1}}(x_{k+1}),x_{k+1})}$ is the DDPM Gaussian approximation of the prior backward transition $q_{{\ell_{k}}|k+1}(\cdot|x_{k+1})$ Ho et al. (2020). In short, we combine the DPS approximation of $g_{{\ell_{k}}}$ and the DDPM approximation of $q_{{\ell_{k}}|k+1}$. Finally, for any sequence $({\ell_{k}})_{k=1}^{n}$ satisfying ${\ell_{k}}\leq k$, we define our surrogate model for $\pi_{0:n}$ and the posterior (2.1) as
$\smash{\hat{\pi}^{\ell}_{0:n}}(x_{0:n})\vcentcolon=\smash{\hat{\pi}_{n}}(x_{n})\prod_{k=0}^{n-1}\smash{\hat{\pi}^{\ell}_{k|k+1}(x_{k}|x_{k+1})}\,,\quad\smash{\hat{\pi}^{\ell}_{0}}(x_{0})\vcentcolon=\int\smash{\hat{\pi}^{\ell}_{0:n}}(x_{0:n})\,\mathrm{d}x_{1:n}\,,$ (3.7)
where $\smash{\hat{\pi}_{n}}(x_{n})=\mathcal{N}(x_{n};\mathrm{0}_{d},\mathrm{I}_{d})$ and $\smash{\hat{\pi}^{\ell}_{0|1}(\cdot|x_{1})}\vcentcolon=\delta_{\smash{m^{\theta}_{0|1}}(x_{1})}$ (similarly to Ho et al., 2020; Song et al., 2021a). Before providing details on the method we propose to sample from $\smash{\hat{\pi}^{\ell}_{0:n}}$, and consequently $\smash{\hat{\pi}^{\ell}_{0}}$, we give an insight into the choice of the sequence $({\ell_{k}})_{k=1}^{n-1}$.
@@
The accuracy of the surrogate model (3.7) ultimately depends on the design of the sequence $({\ell_{k}})_{k=1}^{n-1}$. Indeed, for a given $k$, note that decreasing ${\ell_{k}}$ ensures that the approximation $g\circ\smash{m^{\theta}_{0|{\ell_{k}}}}$ of $g_{{\ell_{k}}}$ becomes more accurate. For instance, setting ${\ell_{k}}=0$
Figure 1: Right: average $W_{2}$ with $10\%$$90\%$ quantile range. Left: distribution of the minimizing $\eta^{*}$.
eliminates any error of the approximate potential (since $g_{0}=g$); however, decreasing ${\ell_{k}}$ can cause the backward transition $\smash{\smash{\pi}_{0|k+1}(\cdot|x_{k+1})}$ to become highly multimodal, especially in the early stages of the diffusion process, making the DDPM Gaussian approximation less accurate. Conversely, setting ${\ell_{k}}$ closer to $k$, similar to the DPS approach of Chung et al. (2023), improves the accuracy of the transition approximation $\smash{q_{{\ell_{k}}|k+1}}$ but worsens the estimate of $\smash{g_{{\ell_{k}}}}$ by $\hat{g}^{\theta}_{{\ell_{k}}}$ as in (3.6). The choice of ${\ell_{k}}$ is thus subject to a particular trade-off that we illustrate using the following Gaussian toy example. When considering the parameterization $\ell^{\eta}_{k}=\lfloor\eta k\rfloor$ with $\eta\in[0,1]$, we show that the Wasserstein-2 distance between $\pi$ and $\smash{\hat{\pi}^{\ell(\eta)}_{0}}$ reaches its minimum w.r.t. $\eta$ for $\eta$ around $0.5$, confirming that the additional flexibility provided by the introduction of a midpoint can provide a surrogate model (3.7) with reduced approximation error.
@@
Consider a linear inverse problem $Y=\mathrm{A}X+\sigma_{y}Z$ with Gaussian noise and prior $q=\mathcal{N}(m,\Sigma)$, where $\mu\in\mathbb{R}^{d}$ and $\Sigma\in\mathcal{S}^{++}_{d}$. In this setting, we have access to the exact denoiser $\smash{m_{0|k}}{}$ in a closed form, which allows us to take $\smash{m^{\theta}_{0|k}}=\smash{m_{0|k}}$. Moreover, since the transition (3.6) admits an explicit expression in this case, we can quantify the approximation error of (3.7) w.r.t. $\pi$ for any sequence $({\ell_{k}})_{k=1}^{n-1}$. Note that in this case, although the true backward transitions $q_{{\ell_{k}}|k+1}$ are Gaussian and have the same mean as the Gaussian DDPM transitions, they differ in their covariance. All necessary derivations and computational details can be found in Appendix B. We consider sequences $(\ell^{\eta}_{k})_{k=1}^{n}$ with $\ell^{\eta}_{k}=\lfloor\eta k\rfloor$, $\eta\in[0,1]$ and compute the Wasserstein-2 distance $W_{2}(\pi,\smash{\hat{\pi}^{\ell(\eta)}_{0}})$ as a function of $\eta$ on the basis of $500$ samples $(\mathrm{A},\mu,\Sigma)$ for ${d}=100$. The left-hand side plot in Figure 1 displays the average error $W_{2}(\pi,\smash{\hat{\pi}^{\ell(\eta)}_{0}})$ as a function of $\eta$ whereas the right-hand side plot displays the associated distribution of $\eta^{*}:=\operatorname*{argmin}_{\eta\in[0,1]}W_{2}(\pi,\smash{\hat{\pi}^{\ell(\eta)}_{0}})$. Figure 1 illustrates that the smallest approximation error for the surrogate model (3.7) is reached at intermediate values close to $\eta=0.5$ rather than $\eta=1$, which corresponds to a DPS-like approximation.
@@
To sample from $\smash{\hat{\pi}^{\ell}_{0:n}}$ in (3.7), we focus on generating approximate samples from the sequence $(\smash{\hat{\pi}^{\ell}_{k}})_{k=0}^{n}$ of marginal distributions of $\smash{\hat{\pi}^{\ell}_{0:n}}$ in (3.5), by backward recursion over $k$. Ideally, for $k=n$, we simply sample from $\smash{\hat{\pi}^{\theta}_{n}}$; then, recursively, assuming that we have access to an approximate sample $X_{k+1}$ from $\smash{\hat{\pi}^{\ell}_{k+1}}$, the next state $X_{k}$ is drawn from $\smash{\hat{\pi}^{\ell}_{k|k+1}(\cdot|X_{k+1})}$. However, as $\smash{\hat{\pi}^{\ell}_{k|k+1}(\cdot|X_{k+1})}$ is intractable due to its definition (3.5), we propose to use the Gaussian approximation that we specify next.
In the following, let $k\in\llbracket 1,n\rrbracket$ and ${\ell_{k}}\in\llbracket 0,k\rrbracket$ be fixed. For $\smash{\varphi=(\hat{\mu}_{{\ell_{k}}},\hat{\rho}_{\ell_{k}})\in(\mathbb{R}^{d})^{2}}$, we consider the Gaussian variational distribution
$\lambda^{\varphi}_{k|k+1}(x_{k}|x_{k+1})$ :$\vcentcolon=\int q_{k|{\ell_{k}},k+1}(x_{k}|x_{\ell_{k}},x_{k+1})\lambda^{\varphi}_{{\ell_{k}}|k+1}(x_{\ell_{k}})\,\mathrm{d}x_{\ell_{k}}\,,$ (3.8)
$\lambda^{\varphi}_{{\ell_{k}}|k+1}(x_{\ell_{k}})$ :$\vcentcolon=\mathcal{N}(x_{\ell_{k}};\hat{\mu}_{\ell_{k}},\mathrm{diag}(\mathrm{e}^{2\hat{\rho}_{\ell_{k}}}))\,.$ (3.9)
In definition (3.9) the exponential is applied element-wise to the vector $\smash{\hat{\rho}_{\ell_{k}}}$ and $\smash{\mathrm{diag}(\mathrm{e}^{\hat{\rho}_{\ell_{k}}})}$ is the diagonal matrix with diagonal entries $\smash{\mathrm{e}^{\hat{\rho}_{\ell_{k}}}}$. Based on the family $\smash{\{\lambda^{\varphi}_{k|k+1}\,:\,\varphi\in(\mathbb{R}^{d})^{2}\}}$ and an approximate sample $X_{k+1}$ from $\smash{\hat{\pi}^{\ell}_{k+1}}$, we then seek the best fitting parameter $\varphi$ that minimizes the upper bound on the backward KL divergence obtained using the data-processing inequality
$\mathsf{KL}(\lambda^{\varphi}_{k|k+1}(\cdot|X_{k+1})\parallel\smash{\hat{\pi}^{\ell}_{k|k+1}(\cdot|X_{k+1})})\leq\mathsf{KL}(\lambda^{\varphi}_{{\ell_{k}}|k+1}\parallel\smash{\hat{\pi}^{\theta}_{{\ell_{k}}|k+1}(\cdot|X_{k+1})})=\vcentcolon\mathcal{L}_{k}(\varphi;X_{k+1})\,.$
@@
$\nabla_{\varphi}\mathcal{L}_{k}(\varphi;X_{k+1})=-\mathbb{E}\big{[}\nabla_{\varphi}\log\hat{g}^{\theta}_{{\ell_{k}}}(\hat{\mu}_{{\ell_{k}}}+\mathrm{diag}(\mathrm{e}^{\hat{\rho}_{\ell_{k}}})Z)\big{]}+\nabla_{\varphi}\mathsf{KL}(\lambda^{\varphi}_{{\ell_{k}}|k+1}\parallel p^{\theta}_{{\ell_{k}}|k+1}(\cdot|X_{k+1}))\,,$
where $Z\sim\mathrm{N}(\mathrm{0}_{d},\mathrm{I}_{d})$ is independent of $X_{k+1}$. The first term is dealt with using the reparameterization trick Kingma & Welling (2014) and can be approximated using a Monte Carlo estimator based on a single sample $Z_{k+1}$; we denote this estimate by $\smash{\nabla_{\varphi}\widetilde{\mathcal{L}}_{k}(\varphi,Z_{k+1};X_{k+1})}$. The second term is the gradient of the KL divergence between two Gaussian distributions and can thus be computed in a closed form. Similarly to the previous approaches of Ho et al. (2022); Chung et al. (2023); Song et al. (2023a), our gradient estimator involves a vector-Jacobian product of the denoising network.
@@
1:  Input: $({\ell_{k}})_{k=1}^{n}$ with ${\ell_{n}}=n$, ${\ell_{1}}=1$, number $M$ of gradient steps.
2:  $X_{n}\sim\mathcal{N}(\mathrm{0}_{d},\mathrm{I}_{d}),\,\hat{X}_{n}\leftarrow X_{n}$

### v2

@@ 3 The MGPS algorithm
We preface our description of the MGPS algorithm with some additional notations. First, consider the DDPM forward process with transitions given by $q_{k|j}(\boldsymbol{x}_{k}|\boldsymbol{x}_{j})\vcentcolon=\mathcal{N}(\boldsymbol{x}_{k};\sqrt{\alpha_{k}/\alpha_{j}}\boldsymbol{x}_{j},(1-\alpha_{k}/\alpha_{j})\mathbf{I}_{d})$ for all $(j,k)\in\llbracket 0,n\rrbracket^{2}$ such that $j<k$. We also define the joint law
$\pi_{0:n}(\boldsymbol{x}_{0:n})\vcentcolon=\pi(\boldsymbol{x}_{0})\prod_{k=0}^{n-1}q_{k+1|k}(\boldsymbol{x}_{k+1}|\boldsymbol{x}_{k})\,,$ (3.1)
of the forward process when initialized with the posterior (2.1) of interest. Note that with these definitions, $\pi_{k}$ (given by (2.3)) is the marginal of $\pi_{0:n}(\boldsymbol{x}_{0:n})$ w.r.t. $\boldsymbol{x}_{k}$. The time-reversed decomposition of (3.1) writes
$\pi_{0:n}(\boldsymbol{x}_{0:n})\vcentcolon=\pi_{n}(\boldsymbol{x}_{n})\prod_{k=0}^{n-1}\smash{\pi}_{k|k+1}(\boldsymbol{x}_{k}|\boldsymbol{x}_{k+1})\,,$ (3.2)
$\smash{\pi}_{i|j}(\boldsymbol{x}_{i}|\boldsymbol{x}_{j})\vcentcolon=\pi_{i}(\boldsymbol{x}_{i})q_{j|i}(\boldsymbol{x}_{j}|\boldsymbol{x}_{i})\big{/}\pi_{j}(\boldsymbol{x}_{j})\propto p_{i}({\boldsymbol{y}}|\boldsymbol{x}_{i})q_{i|j}(\boldsymbol{x}_{i}|\boldsymbol{x}_{j})\,,$ (3.3)
where the marginals are given by (2.2), is the conditional density of $\mathbf{X}_{i}$ given $\mathbf{X}_{j}=\boldsymbol{x}_{j}$ and $\boldsymbol{y}$.
[…]
Figure 1: For each color, the different solid arrows indicate different conditional densities that need to be approximated for a given choice of the midpoint ${\ell_{k}}$. The longer the arrow, the more difficult it is to approximate the corresponding conditional density. By placing the ${\ell_{k}}$ midway between zero and $k$, the shortest arrows are obtained.
Using the decomposition (3.2), an exact draw $\mathbf{X}_{0}$ from (2.1) is obtained by first drawing $\mathbf{X}_{n}$ from $\pi_{n}$ and then simulating recursively $\mathbf{X}_{k}$ from the transitions $\smash{\smash{\pi}_{k|k+1}(\cdot|\mathbf{X}_{k+1})}$ for all $k$. However, according to (3.3), such transitions involve the intractable likelihood $p_{k}({\boldsymbol{y}}|\boldsymbol{x}_{k})$ which is hard to approximate accurately when $k$ is large.
@@
The transition (3.3) has two components: the conditional density $p_{i}({\boldsymbol{y}}|\boldsymbol{x}_{i})$ and the prior transition density $q_{i|j}(\boldsymbol{x}_{i}|\boldsymbol{x}_{j})$. The approximation $p_{i}({\boldsymbol{y}}|\boldsymbol{x}_{i})\approx p({\boldsymbol{y}}|\smash{\boldsymbol{m}^{\theta}_{0|i}}(\boldsymbol{x}_{i}))$ of Chung et al. (2023) is accurate when $i$ is small, whereas the Gaussian approximation of $q_{i|j}(\boldsymbol{x}_{i}|\boldsymbol{x}_{j})$ proposed by Ho et al. (2020) is accurate when the difference $|i-j|$ is small. Thus, in order to combine the best of both worlds, we introduce a decomposition that transfers the problem of sampling from $\smash{\pi}_{k|k+1}$ to that of sampling from $\smash{\pi}_{{\ell_{k}}|k+1}$, i.e., drawing $\mathbf{X}_{\ell_{k}}$ given $\mathbf{X}_{k+1}$ and ${\boldsymbol{y}}$, where ${\ell_{k}}<k$. The parameter ${\ell_{k}}$ balances the approximation errors of the two conditional densities in (3.3); see Figure 1. In order to make this construction, define, for each $\ell\in\llbracket 0,k\rrbracket$, the bridge kernel $q_{k|\ell,k+1}(\boldsymbol{x}_{k}|\boldsymbol{x}_{\ell},\boldsymbol{x}_{k+1})\propto q_{k|\ell}(\boldsymbol{x}_{k}|\boldsymbol{x}_{\ell})q_{k+1|k}(\boldsymbol{x}_{k+1}|\boldsymbol{x}_{k})$, which is Gaussian; see Section A.1. By convention, $q_{k|k,k+1}(\cdot|\boldsymbol{x}_{k},\boldsymbol{x}_{k+1})=\delta_{\boldsymbol{x}_{k}}$.
@@
$\smash{\pi}_{k|k+1}(\boldsymbol{x}_{k}|\boldsymbol{x}_{k+1})=\int q_{k|{\ell_{k}},k+1}(\boldsymbol{x}_{k}|\boldsymbol{x}_{\ell_{k}},\boldsymbol{x}_{k+1})\smash{\pi}_{{\ell_{k}}|k+1}(\boldsymbol{x}_{\ell_{k}}|\boldsymbol{x}_{k+1})\,\mathrm{d}\boldsymbol{x}_{\ell_{k}}\,.$ (3.4)
The proof is found in Section A.3. Lead by the decomposition provided by Lemma 3.1, we define
$\smash{\hat{\pi}^{\ell}_{k|k+1}(\boldsymbol{x}_{k}|\boldsymbol{x}_{k+1})}\vcentcolon=\int q_{k|{\ell_{k}},k+1}(\boldsymbol{x}_{k}|\boldsymbol{x}_{\ell_{k}},\boldsymbol{x}_{k+1})\smash{\hat{\pi}^{\theta}_{{\ell_{k}}|k+1}(\boldsymbol{x}_{\ell_{k}}|\boldsymbol{x}_{k+1})}\,\mathrm{d}\boldsymbol{x}_{{\ell_{k}}}\,,$ (3.5)
where $\smash{\hat{\pi}^{\theta}_{{\ell_{k}}|k+1}(\boldsymbol{x}_{\ell_{k}}|\boldsymbol{x}_{k+1})}\propto\hat{p}^{\theta}_{{\ell_{k}}}({\boldsymbol{y}}|\boldsymbol{x}_{\ell_{k}})p^{\theta}_{{\ell_{k}}|k+1}(\boldsymbol{x}_{\ell_{k}}|\boldsymbol{x}_{k+1})$ with $\hat{p}^{\theta}_{{\ell_{k}}}({\boldsymbol{y}}|\boldsymbol{x}_{\ell_{k}})\vcentcolon=p({\boldsymbol{y}}|\smash{\boldsymbol{m}^{\theta}_{0|{\ell_{k}}}}(\boldsymbol{x}_{\ell_{k}}))$ and $p^{\theta}_{{\ell_{k}}|k+1}(\boldsymbol{x}_{\ell_{k}}|\boldsymbol{x}_{k+1})\vcentcolon=q_{{\ell_{k}}|0,k+1}(\boldsymbol{x}_{\ell_{k}}|\smash{\boldsymbol{m}^{\theta}_{0|k+1}}(\boldsymbol{x}_{k+1}),\boldsymbol{x}_{k+1})$. Finally, for any sequence :$\ell\vcentcolon=({\ell_{k}})_{k=1}^{n}$ satisfying ${\ell_{k}}\leq k$, we define our surrogate model for $\pi_{0:n}$ and the posterior (2.1) as
$\smash{\hat{\pi}^{\ell}_{0:n}}(\boldsymbol{x}_{0:n})\vcentcolon=\smash{\hat{\pi}_{n}}(\boldsymbol{x}_{n})\prod_{k=0}^{n-1}\smash{\hat{\pi}^{\ell}_{k|k+1}(\boldsymbol{x}_{k}|\boldsymbol{x}_{k+1})}\,,\quad\smash{\hat{\pi}^{\ell}_{0}}(\boldsymbol{x}_{0})\vcentcolon=\int\smash{\hat{\pi}^{\ell}_{0:n}}(\boldsymbol{x}_{0:n})\,\mathrm{d}\boldsymbol{x}_{1:n}\,,$ (3.6)
where $\smash{\hat{\pi}_{n}}(\boldsymbol{x}_{n})=\mathcal{N}(\boldsymbol{x}_{n};\mathbf{0}_{d},\mathbf{I}_{d})$ and $\smash{\hat{\pi}^{\ell}_{0|1}(\cdot|\boldsymbol{x}_{1})}\vcentcolon=\delta_{\smash{\boldsymbol{m}^{\theta}_{0|1}}(\boldsymbol{x}_{1})}$ (similarly to Ho et al., 2020; Song et al., 2021a). Since the bridge kernel $q_{k|{\ell_{k}},k+1}$ is Gaussian, it can be easily sampled. Hence, we only need to focus on the approximate sampling from $\smash{\smash{\pi}_{{\ell_{k}}|k+1}}$. Moreover, since the decomposition (3.4) is valid for any ${\ell_{k}}$, it can be chosen to better balance the approximation errors, as discussed above.
@@
The accuracy of the surrogate model (3.6) ultimately depends on the design of the sequence $({\ell_{k}})_{k=1}^{n-1}$. In fact, for a given $k$, note that a decrease in ${\ell_{k}}$ ensures that the approximation $\hat{p}^{\theta}_{{\ell_{k}}}({\boldsymbol{y}}|\boldsymbol{x}_{\ell_{k}})$ of $p_{{\ell_{k}}}({\boldsymbol{y}}|\boldsymbol{x}_{\ell_{k}})$ becomes more accurate. For instance, setting ${\ell_{k}}=0$
Figure 2: Left: average $W_{2}$ with $10\%$$90\%$ quantile range. Right: distribution of the minimizing $\eta^{*}$.
eliminates any error of the approximate likelihood (since $p_{0}({\boldsymbol{y}}|\cdot)=p({\boldsymbol{y}}|\cdot)$); however, decreasing ${\ell_{k}}$ can cause the distribution $\smash{\smash{\pi}_{0|k+1}(\cdot|\boldsymbol{x}_{k+1})}$ to become strongly multimodal, especially in the early stages of the diffusion process, making the DDPM Gaussian approximation less accurate. Conversely, setting ${\ell_{k}}$ closer to $k$, similar to the DPS approach of Chung et al. (2023), improves the accuracy of the approximation of $\smash{q_{{\ell_{k}}|k+1}}$, but impairs the approximation $\hat{p}^{\theta}_{{\ell_{k}}}({\boldsymbol{y}}|\boldsymbol{x}_{\ell_{k}})$ of $\smash{p_{{\ell_{k}}}({\boldsymbol{y}}|\boldsymbol{x}_{\ell_{k}})}$. The choice of ${\ell_{k}}$ is therefore subject to a particular trade-off, which we illustrate using the following Gaussian toy example. When considering the parameterization $\ell^{\eta}_{k}=\lfloor\eta k\rfloor$ with $\eta\in[0,1]$, we show that the Wasserstein-2 distance between $\pi$ and $\smash{\hat{\pi}^{\ell(\eta)}_{0}}$ reaches its minimum for $\eta$ around $0.5$, which confirms that the additional flexibility obtained by introducing the midpoint can provide a surrogate model (3.6) with reduced approximation error.
@@
Consider a linear inverse problem $\mathbf{Y}=\mathbf{A}\mathbf{X}+\sigma_{\boldsymbol{y}}\mathbf{Z}$ with Gaussian noise and prior $q=\mathcal{N}(\boldsymbol{m},\mathbf{\Sigma})$, where $\boldsymbol{m}\in\mathbb{R}^{d}$ and $\mathbf{\Sigma}\in\mathbb{R}^{d\times d}$ is positive definite. In this setting, we have access to the exact denoiser $\smash{\boldsymbol{m}_{0|k}}{}$ in a closed form, which allows us to take $\smash{\boldsymbol{m}^{\theta}_{0|k}}=\smash{\boldsymbol{m}_{0|k}}$. Moreover, since the transition (3.5) admits an explicit expression in this case, we can quantify the approximation error of (3.6) w.r.t. $\pi$ for any sequence $({\ell_{k}})_{k=1}^{n-1}$. Note that in this case the true backward transitions $q_{{\ell_{k}}|k+1}$ are Gaussian and have the same mean as the Gaussian DDPM transitions, but differ in their covariance. All necessary derivations and computational details can be found in Appendix B. We consider sequences $\ell(\eta)\vcentcolon=(\ell^{\eta}_{k})_{k=1}^{n}$ with $\ell^{\eta}_{k}=\lfloor\eta k\rfloor$, $\eta\in[0,1]$, and compute the Wasserstein-2 distance $W_{2}(\pi,\smash{\hat{\pi}^{\ell(\eta)}_{0}})$ as a function of $\eta$ on the basis of $500$ samples $(\mathbf{A},\boldsymbol{m},\mathbf{\Sigma})$ for ${d}=100$. The left panel of Figure 2 displays the average error $W_{2}(\pi,\smash{\hat{\pi}^{\ell(\eta)}_{0}})$ as a function of $\eta$, while the right panel displays the associated distribution of $\eta^{*}\vcentcolon=argmin_{\eta\in[0,1]}W_{2}(\pi,\smash{\hat{\pi}^{\ell(\eta)}_{0}})$. Figure 2 illustrates that the smallest approximation error for the surrogate model (3.6) is reached at intermediate values close to $\eta=0.5$ rather than $\eta=1$, which corresponds to a DPS-like approximation.
@@
To sample from $\smash{\hat{\pi}^{\ell}_{0:n}}$ in (3.6), we focus on simulating approximately and recursively (over $k$) the Markov chain with transition densities (3.5). Ideally, for $k=n$, we simply sample from $\smash{\hat{\pi}^{\theta}_{n}}$; then, recursively, assuming that we have access to an approximate sample $\mathbf{X}_{k+1}$ from $\smash{\hat{\pi}^{\ell}_{k+1}}$, the next state $\mathbf{X}_{k}$ is drawn from $\smash{\hat{\pi}^{\ell}_{k|k+1}(\cdot|\mathbf{X}_{k+1})}$. However, as $\smash{\hat{\pi}^{\ell}_{k|k+1}(\cdot|\mathbf{X}_{k+1})}$ is intractable, we propose using the Gaussian approximation that we specify next.
In the following, let $k\in\llbracket 1,n\rrbracket$ and ${\ell_{k}}\in\llbracket 0,k\rrbracket$ be fixed. For $\smash{{\boldsymbol{\varphi}}=(\hat{\boldsymbol{\mu}}_{{\ell_{k}}},\hat{\boldsymbol{\rho}}_{\ell_{k}})\in(\mathbb{R}^{d})^{2}}$, consider the Gaussian variational distribution
$\lambda^{{\boldsymbol{\varphi}}}_{k|k+1}(\boldsymbol{x}_{k}|\boldsymbol{x}_{k+1})$ :$\vcentcolon=\int q_{k|{\ell_{k}},k+1}(\boldsymbol{x}_{k}|\boldsymbol{x}_{\ell_{k}},\boldsymbol{x}_{k+1})\lambda^{{\boldsymbol{\varphi}}}_{{\ell_{k}}|k+1}(\boldsymbol{x}_{\ell_{k}})\,\mathrm{d}\boldsymbol{x}_{\ell_{k}}\,,$ (3.7)
$\lambda^{{\boldsymbol{\varphi}}}_{{\ell_{k}}|k+1}(\boldsymbol{x}_{\ell_{k}})$ :$\vcentcolon=\mathcal{N}(\boldsymbol{x}_{\ell_{k}};\hat{\boldsymbol{\mu}}_{\ell_{k}},\mbox{diag}(\mathrm{e}^{2\hat{\boldsymbol{\rho}}_{\ell_{k}}}))\,.$ (3.8)
In definition (3.8) the exponential function is applied element-wise to the vector $\smash{\hat{\boldsymbol{\rho}}_{\ell_{k}}}$ and $\smash{\mbox{diag}(\mathrm{e}^{\hat{\boldsymbol{\rho}}_{\ell_{k}}})}$ is the diagonal matrix with diagonal entries $\smash{\mathrm{e}^{\hat{\boldsymbol{\rho}}_{\ell_{k}}}}$. Based on the family $\smash{\{\lambda^{{\boldsymbol{\varphi}}}_{k|k+1}\,:\,{\boldsymbol{\varphi}}\in(\mathbb{R}^{d})^{2}\}}$ and an approximate sample $\mathbf{X}_{k+1}$ from $\smash{\hat{\pi}^{\ell}_{k+1}}$, we then seek the best fitting parameter ${\boldsymbol{\varphi}}$ that minimizes the upper bound on the backward KL divergence obtained using the data-processing inequality
$\mathsf{KL}(\lambda^{{\boldsymbol{\varphi}}}_{k|k+1}(\cdot|\mathbf{X}_{k+1})\parallel\smash{\hat{\pi}^{\ell}_{k|k+1}(\cdot|\mathbf{X}_{k+1})})\leq\mathsf{KL}(\lambda^{{\boldsymbol{\varphi}}}_{{\ell_{k}}|k+1}\parallel\smash{\hat{\pi}^{\theta}_{{\ell_{k}}|k+1}(\cdot|\mathbf{X}_{k+1})})=\vcentcolon\mathcal{L}_{k}({\boldsymbol{\varphi}};\mathbf{X}_{k+1})\,.$
@@
$\nabla_{\boldsymbol{\varphi}}\mathcal{L}_{k}({\boldsymbol{\varphi}};\mathbf{X}_{k+1})=-\mathbb{E}\big{[}\nabla_{\boldsymbol{\varphi}}\log\hat{p}^{\theta}_{{\ell_{k}}}({\boldsymbol{y}}|\hat{\boldsymbol{\mu}}_{{\ell_{k}}}+\mbox{diag}(\mathrm{e}^{\hat{\boldsymbol{\rho}}_{\ell_{k}}})\mathbf{Z})\big{]}+\nabla_{\boldsymbol{\varphi}}\mathsf{KL}(\lambda^{{\boldsymbol{\varphi}}}_{{\ell_{k}}|k+1}\parallel p^{\theta}_{{\ell_{k}}|k+1}(\cdot|\mathbf{X}_{k+1}))\,,$
where $\mathbf{Z}\sim\mathrm{N}(\mathbf{0}_{d},\mathbf{I}_{d})$ is independent of $\mathbf{X}_{k+1}$. The first term is dealt with using the reparameterization trick Kingma & Welling (2014) and can be approximated using a Monte Carlo estimator based on a single sample $\mathbf{Z}_{k+1}$; we denote this estimate by $\smash{\nabla_{\boldsymbol{\varphi}}\widetilde{\mathcal{L}}_{k}({\boldsymbol{\varphi}},\mathbf{Z}_{k+1};\mathbf{X}_{k+1})}$. The second term is the gradient of the KL divergence between two Gaussian distributions and can thus be computed in a closed form. Similarly to the previous approaches of Ho et al. (2022); Chung et al. (2023); Song et al. (2023a), our gradient estimator involves a vector-Jacobian product of the denoising network.
@@
1:  Input: $({\ell_{k}})_{k=1}^{n}$ with ${\ell_{n}}=n$ and ${\ell_{1}}=1$; number $M$ of gradient steps.
2:  $\mathbf{X}_{n}\sim\mathcal{N}(\mathbf{0}_{d},\mathbf{I}_{d}),\,\hat{\mathbf{X}}_{n}\leftarrow\mathbf{X}_{n}$